OpenStudy (anonymous):

Determine the following limits, if it exists: 3) lim x->0 (cube rootx)-1 / (square rootx)-1 4) lim n->infinity n(n)!/(n+2)!

OpenStudy (freckles):

(n+2)!=(n+2)(n+1)*n!

OpenStudy (freckles):

so you should be able to write 4 without factorials

OpenStudy (freckles):

3 says \[\lim_{x \rightarrow 0}\frac{\sqrt[3]{x}-1}{\sqrt{x}-1}\]?

OpenStudy (freckles):

if so try a sub

OpenStudy (freckles):

\[\text{ \to get rid of the ugly radicals \choose a sub such that its power } \\ \text{ is divisible by both 3 and 2 }\] like 6 so we choose the following sub: \[x=u^6 \]

OpenStudy (freckles):

solving that for u we see as x goes to 0 u also goes to 0

OpenStudy (freckles):

\[\lim_{u \rightarrow 0} \frac{u^\frac{6}{3}-1}{x^\frac{6}{2}-1}\] say by by the ugly radicals

OpenStudy (freckles):

now you reduce the powers and factor

OpenStudy (freckles):

and cancel

OpenStudy (freckles):

and plug in

OpenStudy (anonymous):

can't you rationalize denominator then factor a difference of cubes

OpenStudy (freckles):

\[(\sqrt[3]{x}-1)(\sqrt{x}+1)=x^{\frac{1}{3}+\frac{1}{2}}+x^\frac{1}{3}-x^\frac{1}{2}-1 =x^\frac{5}{6}+x^\frac{1}{3}-x^\frac{1}{2}-1\] if i'm not mistaken this is what you get for the numerator when trying to rationlize the denominator as is

OpenStudy (freckles):

I think the sub simplifies things

OpenStudy (freckles):

\[\lim_{u \rightarrow 0} \frac{u^\frac{6}{3}-1}{u^\frac{6}{2}-1}=\lim_{u \rightarrow 0}\frac{u^2-1}{u^3-1}=\lim_{u \rightarrow 0}\frac{(u-1)(u+1)}{(u-1)(u^2+u+1)}\]

OpenStudy (anonymous):

so it will be 1?

OpenStudy (freckles):

yea (0+1)/(0^2+0+1)=1/1=1

OpenStudy (anonymous):

thanks lots! but i dont understand the 4) question.

OpenStudy (freckles):

do you understand that (n+2)!=(n+2)*(n+1)*n!

OpenStudy (anonymous):

no

OpenStudy (freckles):

do you know what 5! equals?

OpenStudy (anonymous):

120?

OpenStudy (freckles):

and how did you get that

OpenStudy (anonymous):

graphing calculator...

OpenStudy (freckles):

I was looking for this \[5!=5 \cdot 4 \cdot 3 \cdot 2 \cdot 1 \]

OpenStudy (anonymous):

oh yeah...

OpenStudy (freckles):

\[n!=n \cdot (n-1) \cdot (n-2) \cdot (n-3) \cdots 3 \cdot 2 \cdot 1 \]

OpenStudy (anonymous):

okay...

OpenStudy (freckles):

\[5!=5 \cdot 4 \cdot 3 \cdot 2 \cdot 1 \text{ or we can write \it as } \\ 5!=5 \cdot 4! \text{ or even } 5!=5 \cdot 4 \cdot 3! \]

OpenStudy (freckles):

\[(n+2)!=(n+2) \cdot (n+1) \cdot n! \] I stopped at the n part here because I have n! on top and bottom now

OpenStudy (freckles):

\[\lim_{n \rightarrow \infty}\frac{n \cdot n!}{(n+2)(n+1) \cdot n!}\]

OpenStudy (freckles):

\[\lim_{n \rightarrow \infty}\frac{n}{(n+2)(n+1)}\]

OpenStudy (anonymous):

oh i see.. how would i sub in the infinity?

OpenStudy (freckles):

multiply the bottom out and use that \[\lim_{x \rightarrow \pm \infty}\frac{1}{x^r}=0 \text{ where } r>0\]

OpenStudy (freckles):

\[\lim_{n \rightarrow \infty}\frac{n}{n^2+3n+2} =\lim_{n \rightarrow \infty}\frac{\frac{n}{n^2}}{\frac{n^2}{n^2}+\frac{3n}{n^2}+\frac{2}{n^2}}\]

OpenStudy (freckles):

I must go now but you should be able to do this part now

OpenStudy (anonymous):

thank you sooo much!